OpenStudy (tukitw):

The function f is defined by f(x) = x^2 - kx + 3, where the domain of f is the set of all real numbers and k is a non-zero constant. Given that the range of f is [-1 , infinity) and that f^-1 exists if the domain of f is restricted to the set of negative real numbers, find the value of k.

OpenStudy (anonymous):

\[-1\le x^2-kx+3< \infty \]

OpenStudy (anonymous):

we don't have any idea about infinity so we will go with \[-1 \le x^2-kx+3\]

OpenStudy (anonymous):

x^2-kx+3+1>=0 x^2-kx+4>=0

OpenStudy (anonymous):

it is quadratic equation

OpenStudy (anonymous):

so for it to be >=0 D>=0 k^2-4*1*4>=0 k^2-16>=0 (k-4)(k+4)>=0

OpenStudy (anonymous):

k>=4 or k<=-4

OpenStudy (tukitw):

The answer given is only k = 4. Any idea why?

OpenStudy (tukitw):

Nevermind, I managed to solve it. Since f(x) is a positive quadratic curve, it has a minimum point. f(x) = x^2 - kx + 3 f'(x) = 2x - k For f(x) to be minimum, f'(x) = 0. 2x - k = 0 x = k / 2 Since minimum f = -1, f(k / 2) = -1 (k / 2)^2 - k(k / 2) + 3 = -1 k^2 / 4 - k^2 / 2 + 4 = 0 k^2 - 2k^2 + 16 = 0 k^2 - 16 = 0 k = 4 or k = -4 Next, the domain of f is restricted to (-inf , 0). Assume k = -4, f'(x) = 2x + 4, x < 0 Since f'(x) changes from positive to negative, f(x) is not strictly increasing or decreasing as x increases. Thus, f^-1 does not exists, which contradicts the question. Assume k = 4, f'(x) = 2x - 4, x < 0 Since f'(x) remains negative, f(x) is strictly decreasing as x increases. Thus, f^-1 exists. So, k = 4.